# DeFoG: Discrete Flow Matching for Graph Generation

### v1

@@ C.2 Graph Specific Theoretical Results
C.2.1 Additional Features Expressivity
@@
Siraudin et al. (2024) experimentally validate the effectiveness of RRWP features for graph diffusion models and propose extending their proof to additional graph properties that GNNs fail to capture (Xu et al., 2019; Morris et al., 2018). For example, an MLP with input $M^{k}$ with $k=N-1$ for an $N$-node graph can approximate the connected component of each node and the number of vertices in the largest connected component. Additionally, RRWP features can be used to capture cycle-related information.
C.2.2 Node Permutation Equivariance and Invariance Properties
The DeFoG model is permutation equivariant, its loss function is permutation invariant, and its sampling probability is permutation invariant.
Proof.
Recall that we denote an undirected graph with $N$ nodes by $G=(x^{1:n:N},e^{1:i<j:N})$. Here, each node variable is represented as $x^{n}\in\mathcal{X}=\{1,\dots,X\}$, and each edge variable as $e^{ij}\in\mathcal{E}=\{1,\dots,E\}$. We also treat $G$ as a multivariate data point consisting of $D$ discrete variables including all nodes and all edges.
Permutation Equivariant Model
We begin by proving that the DeFoG model is permutation-equivariant, including the network architecture and the additional features employed.
[…]
    We first prove that $M(A)=D^{-1}A$ is permutation equivariant:
    ${M(A^{\prime})}^{ij}$ $=(D^{\prime-1}A^{\prime})^{ij}$
    $=(1/(D^{\prime})^{ii})(A^{\prime})^{ij}$
    $=(D^{-1}A)^{(\sigma^{-1}(i),\sigma^{-1}(j))}$
    $={(M(A)^{\prime})}^{ij}.$
    ${P(\pi M\pi^{T})}$ $=[\pi I\pi^{T},\pi M\pi^{T},\dots,(\pi M\pi^{T})^{K-1}]$
    $=\pi[I,M,\dots,M^{K-1}]\pi^{T}$
    $={\pi P(M)\pi^{T}}.$
Hence, since all of its components are permutation-equivariant, so is the DeFoG full architecture.
Permutation Invariant Loss Function
Permutation Invariant Sampling Probability
$=$ $R^{*,n}_{t}\left(x^{\prime(n)}_{t},x^{\prime(n)}_{t+\mathrm{d}t}|x^{\prime(n)}_{1}\right)$
$=$ $R^{\sigma^{-1}(n)}_{t}(x^{\sigma^{-1}(n)}_{t},x^{\sigma^{-1}(n)}_{t+\mathrm{d}t}|x^{\sigma^{-1}(n)}_{1})$
$\tilde{p}_{t+\Delta t|t}(G^{1:D}_{t+\Delta t}|G^{1:D}_{t})$ $=$ $\prod_{n}$ $\tilde{p}^{n}_{t+\Delta t|t}(x^{n}_{t+\Delta t}|G_{t})\prod_{i<j}\tilde{p}^{ij}_{t+\Delta t|t}(e^{ij}_{t+\Delta t}|G_{t})$
$=$ $\prod_{n}$ $\left(\delta(x^{n}_{t},x^{n}_{t+\Delta t})+\mathbb{E}_{p^{\theta,n}_{1|t}(x^{n}_{1}|G_{t})}\left[R^{n}_{t}(x^{n}_{t},x^{n}_{t+\Delta t}|x^{n}_{1})\right]\,\Delta t\right)$
$\prod_{i<j}\left(\delta(e^{ij}_{t},e^{ij}_{t+\Delta t})+\mathbb{E}_{p^{\theta,ij}_{1|t}(e^{ij}_{1}|G_{t})}\left[R^{ij}_{t}(e^{ij}_{t},e^{ij}_{t+\Delta t}|e^{ij}_{1})\right]\,\Delta t\right)$
The transition probability $\tilde{p}_{t+\Delta t|t}(G_{t+\Delta t}|G_{t})$ is expressed as a product over all nodes and edges, an operation that is inherently a permutation invariant function with respect to node ordering. Furthermore, as demonstrated earlier, the terms $\mathbb{E}_{p^{\theta,n}_{1|t}(x^{n}_{1}|G_{t})}\left[R^{n}_{t}(x^{n}_{t},x^{n}_{t+\Delta t}|x^{n}_{1})\right]$ and $\mathbb{E}_{p^{\theta,ij}_{1|t}(e^{ij}_{1}|G_{t})}\left[R^{ij}_{t}(e^{ij}_{t},e^{ij}_{t+\Delta t}|e^{ij}_{1})\right]$ are permutation equivariant since $R^{n}_{t}$, $p^{\theta,n}_{1|t}(x^{n}_{1}|G_{t})$, $R^{ij}_{t}$, and $p^{\theta,ij}_{1|t}(e^{ij}_{1}|G_{t})$ are permutation equivariant if model $f_{\theta}$ is permutation equivariant. Consequently, since the composition of these components yields a permutation invariant function, we conclude that the transition probability of the considered CTMC is permutation invariant.
We finally verify the final sampling probability, $\tilde{p}_{1}(G^{1:D}_{1})$, is permutation-invariant. To simulate $G_{1}^{1:D}$ over $K$ time steps $[0=t_{0},t_{1},\ldots,t_{K}=1]$, we can marginalize it first by taking the expectation over the state at the last time step $t_{K-1}$. Specifically, we have $p_{1}(G^{1:D}_{1})=\mathbb{E}_{p(G^{1:D}_{t_{K-1}})}\left[\tilde{p}_{1|t}(G^{1:D}_{1}|G^{1:D}_{t_{K-1}})\right]$. Since the process is Markovian, this expression can be sequentially extended over the $T$ steps through successive expectations.
$p_{1}(G^{1:D}_{1})$ $=\mathbb{E}_{p(G^{1:D}_{t_{K-1}})}\left[\tilde{p}_{1|t_{K-1}}(G^{1:D}_{1}|G^{1:D}_{t_{K-1}})\right]$
$=\mathbb{E}_{p(G^{1:D}_{t_{K-2}})}\underbrace{\left[\mathbb{E}_{\tilde{p}_{t_{K-1}|t_{K-2}}(G^{1:D}_{t_{K-1}}|G^{1:D}_{t_{K-2}})}\left[\tilde{p}_{1|t_{K-1}}(G^{1:D}_{1}|G^{1:D}_{t_{K-1}})\right]\right]}_{\tilde{p}_{1|t_{K-2}}(G^{1:D}_{1}|G^{1:D}_{t_{K-2}})}$
$\ldots$
$=\mathbb{E}_{p(G^{1:D}_{0})}\left[\mathbb{E}_{\tilde{p}_{t_{1}|0}(G^{1:D}_{t_{1}}|G^{1:D}_{0})}\left[\ldots\mathbb{E}_{\tilde{p}_{t_{K-1}|t_{K-2}}(G^{1:D}_{t_{K-1}}|G^{1:D}_{t_{K-2}})}\left[\tilde{p}_{1|t_{K-1}}(G^{1:D}_{1}|G^{1:D}_{t_{K-1}})\right]\right]\right]$
Due to the fact that each function in the sequence is itself permutation-invariant, the composition of permutation-invariant functions preserves this invariance throughout. Thus, the final sampling probability is invariant over isomorphic graphs. ∎
Appendix D Conditional Generation
@@
where $\gamma$ denotes the guidance weight. In particular, the case with $\gamma=1$ corresponds to standard conditional generation, while $\gamma=0$ represents standard unconditional generation. As $\gamma$ increases, the conditioning effect described by ${\left(\frac{R^{\theta}_{t}(x,\tilde{x}|y)}{R^{\theta}_{t}(x,\tilde{x})}\right)}^{\gamma-1}$ is strengthened, thereby enhancing the quality of the generated samples. We observed $\gamma=2.0$ to be the best performing value for our digital pathology experiments (Sec. 6.4), as detailed in Tab. 6.
@@
Appendix E Experimental Details
@@
E.1 Dataset Details
E.1.1 Synthetic Datasets
Here, we describe the datasets employed in our experiments and outline the specific metrics used to evaluate model performance on each dataset. Additional visualizations of example graphs from each dataset, along with generated graphs, are provided in Figures 21, 23 and 25.
@@ E.1.2 Molecular datasets
Finally, we include the Guacamol benchmark (Brown et al., 2019), based on the ChEMBL 24 database (Mendez et al., 2019). This dataset comprises synthesized molecules, tested against biological targets, with sizes ranging from 2 to 88 heavy atoms.
Next, we evaluate DeFoG on the Moses benchmark (Polykovskiy et al., 2020), derived from the ZINC Clean Leads collection (Sterling & Irwin, 2015), which contains molecules ranging from 8 to 27 heavy atoms, filtered according to certain criteria. Lastly, we include the Guacamol benchmark (Brown et al., 2019), derived from the ChEMBL 24 database (Mendez et al., 2019). This benchmark contains only molecules that have been synthesized and tested against biological targets, with 2 to 88 heavy atoms.
@@
For the QM9 dataset, we follow the dataset splits and evaluation metrics outlined by Vignac et al. (2022). For the Moses and Guacamol benchmarks, we adhere to the training setups and evaluation metrics proposed by Polykovskiy et al. (2020) and Brown et al. (2019), respectively. Note that Guacamol includes molecules with charges; therefore, the generated graphs are converted to charged molecules based on the relaxed validity criterion used by Jo et al. (2022) before being translated to their corresponding SMILES representations. The validity, uniqueness, and novelty metrics reported by the Guacamol benchmark are actually V, V.U., and V.U.N., and are referred to directly as V, V.U., and V.U.N. in the table for clarity.
E.1.3 Digital Pathology Datasets
@@
Graphs, with their natural ability to represent relational data, are widely used to capture spatial biological dependencies in tissue images. This approach has proven successful in digital pathology tasks such as microenvironment classification (Wu et al., 2022), cancer classification (Pati et al., 2022), and decision explainability (Jaume et al., 2021). More recently, graph-based methods have been applied to generative tasks (Madeira et al., 2023), and an open-source dataset was made available by Madeira et al. (2024). This dataset consists of cell graphs where the nodes represent biological cells, categorized into 9 distinct cell types (node classes), and edges model local cell-cell interactions (a single class). For further details, refer to Madeira et al. (2024).
@@
Each cell graph in the dataset can be mapped to a TLS (Tertiary Lymphoid Structure) embedding, denoted as $\kappa=[\kappa_{0},\ldots,\kappa_{5}]\in\mathbb{R}^{6}$, which quantifies its TLS content. A graph $G$ is classified as having low TLS content if $\kappa_{1}(G)<0.05$, and high TLS content if $\kappa_{2}(G)>0.05$. Based on these criteria, the dataset is split into two subsets: high TLS and low TLS. In prior work, TLS generation accuracy was evaluated by training generative models on these subsets separately, and verifying if the generated graphs matched the corresponding TLS content label. We compute TLS accuracy as the average accuracy across both subsets. For DeFoG, we conditionally train it on both subsets simultaneously, as described in Appendix D, and compute TLS accuracy based on whether the generated graphs adhere to the conditioning label. Additionally, we report the V.U.N. metric (valid, unique, novel), similar to what is done for the synthetic datasets (see Sec. E.1.1). A graph is considered valid in this case if it is a connected planar graph, as the graphs in these datasets were constructed using Delaunay triangulation.
E.2 Resources
@@
E.3 Hyperparameter Tuning
The default hyperparameters for training and sampling for each dataset can be found in the provided code repository. In Tab. 6, we specifically highlight their values for the proposed training (see Sec. 4.2.1) and sampling (see Sec. 4.2.2) strategies, and conditional guidance parameter (see Appendix D). As the training process is by far the most computationally costly stage, we aim to minimize changes to the default model training configuration. Nevertheless, we demonstrate the effectiveness of these modifications on certain datasets:

### v3

@@ D.2 Theoretical Results on Architectural Expressivity
D.2.1 Node Permutation Equivariance and Invariance Properties
For any permutation-equivariant denoising neural network, the loss function of DeFoG is permutation invariant, and its sampling probability is permutation invariant.
Proof.
Recall that we denote an undirected graph with $N$ nodes by $G=(x^{1:n:N},e^{1:i<j:N})$. Here, each node variable is represented as $x^{n}\in\mathcal{X}=\{1,\dots,X\}$, and each edge variable as $e^{(ij)}\in\mathcal{E}=\{1,\dots,E\}$. We also treat $G$ as a multivariate data point consisting of $D$ discrete variables including all nodes and all edges.
Permutation Equivariant Model
We begin by proving that the DeFoG architecture is permutation-equivariant, including the network architecture and the additional features employed.
[…]
    We first prove that $M(A)=D^{-1}A$ is permutation equivariant:
    ${M(A^{\prime})}^{(ij)}$ $=(D^{\prime-1}A^{\prime})^{(ij)}$
    $=(1/(D^{\prime})^{(ii)})(A^{\prime})^{(ij)}$
    $=(D^{-1}A)^{(\sigma^{-1}(i),\sigma^{-1}(j))}$
    $={(M(A)^{\prime})}^{(ij)}.$
    ${P(\pi M\pi^{T})}$ $=[\pi I\pi^{T},\pi M\pi^{T},\dots,(\pi M\pi^{T})^{K-1}]$
    $=\pi[I,M,\dots,M^{K-1}]\pi^{T}$
    $={\pi P(M)\pi^{T}}.$
Hence, since all of its components are permutation-equivariant, so is the DeFoG full architecture. The next results hold for any permutation equivariant denoising neural network.
Permutation Invariant Loss Function
Permutation Invariant Sampling Probability
$=$ $R^{*}_{t}\left(x^{\prime(n)}_{t},x^{\prime(n)}_{t+\mathrm{d}t}|x^{\prime(n)}_{1}\right)$
$=$ $R^{*}_{t}(x^{\sigma^{-1}(n)}_{t},x^{\sigma^{-1}(n)}_{t+\mathrm{d}t}|x^{\sigma^{-1}(n)}_{1})$
$\tilde{p}_{t+\Delta t|t}(G^{1:D}_{t+\Delta t}|G^{1:D}_{t})$ $=\prod_{d=1}^{D}\tilde{p}^{(d)}_{t+\Delta t|t}(G^{(d)}_{t+\Delta t}|G^{1:D}_{t})$
$=\prod_{d=1}^{D}\left(\delta(G^{(d)}_{t},G^{(d)}_{t+\Delta t})+\mathbb{E}_{p^{\theta}(G^{(d)}_{1}|G^{1:D}_{t})}\left[R^{(d)}_{t}(G^{(d)}_{t},G^{(d)}_{t+\Delta t}|G^{(d)}_{1})\right]\,\Delta t\right).$
The transition probability $\tilde{p}_{t+\Delta t|t}(G^{1:D}_{t+\Delta t}|G^{1:D}_{t})$ is expressed as a product over all nodes and edges, an operation that is inherently a permutation invariant function with respect to node ordering. Furthermore, as demonstrated earlier, the term $\mathbb{E}_{p^{\theta}(G^{(d)}_{1}|G^{1:D}t)}\left[R^{(d)}_{t}(G^{(d)}_{t},G^{(d)}_{t+\Delta t}|G^{(d)}_{1})\right]$ is permutation equivariant since $R^{(d)}_{t}$ and $p_{t|1}^{\theta}(G^{(d)}_{1}|G^{1:D}_{t})$ are both permutation equivariant if model $f_{\theta}$ is permutation equivariant. Consequently, since the composition of these components yields a permutation invariant function, we conclude that the transition probability of the considered CTMC is permutation invariant.
We finally verify the final sampling probability, $\tilde{p}_{1}(G^{1:D}_{1})$, is permutation-invariant. To simulate $G_{1}^{1:D}$ over $K$ time steps $[0=t_{0},t_{1},\ldots,t_{K}=1]$, we can marginalize it first by taking the expectation over the state at the last time step $t_{K-1}$. Specifically, we have $p_{1}(G^{1:D}_{1})=\mathbb{E}_{p_{t_{K-1}}(G^{1:D}_{t_{K-1}})}\left[\tilde{p}_{1|t_{K-1}}(G^{1:D}_{1}|G^{1:D}_{t_{K-1}})\right]$. Since the process is Markovian, this expression can be sequentially extended over the $T$ steps through successive expectations.
$p_{1}(G^{1:D}_{1})$ $=\mathbb{E}_{p_{t_{K-1}}(G^{1:D}_{t_{K-1}})}\left[\tilde{p}_{1|t_{K-1}}(G^{1:D}_{1}|G^{1:D}_{t_{K-1}})\right]$
$=\mathbb{E}_{p_{t_{K-2}}(G^{1:D}_{t_{K-2}})}\underbrace{\left[\mathbb{E}_{\tilde{p}_{t_{K-1}|t_{K-2}}(G^{1:D}_{t_{K-1}}|G^{1:D}_{t_{K-2}})}\left[\tilde{p}_{1|t_{K-1}}(G^{1:D}_{1}|G^{1:D}_{t_{K-1}})\right]\right]}_{\tilde{p}_{1|t_{K-2}}(G^{1:D}_{1}|G^{1:D}_{t_{K-2}})}$
$\ldots$
$=\mathbb{E}_{p_{0}(G^{1:D}_{0})}\left[\mathbb{E}_{\tilde{p}_{t_{1}|0}(G^{1:D}_{t_{1}}|G^{1:D}_{0})}\left[\ldots\mathbb{E}_{\tilde{p}_{t_{K-1}|t_{K-2}}(G^{1:D}_{t_{K-1}}|G^{1:D}_{t_{K-2}})}\left[\tilde{p}_{1|t_{K-1}}(G^{1:D}_{1}|G^{1:D}_{t_{K-1}})\right]\right]\right]$
Due to the fact that each function in the sequence is itself permutation-invariant and that the initial distribution $p_{0}(G^{1:D}_{0})$ is permutation invariant (see Sec. C.1), the composition of permutation-invariant functions preserves this invariance throughout. Thus, the final sampling probability is invariant over isomorphic graphs. ∎
D.2.2 Additional Features Expressivity
@@
Siraudin et al. (2024) experimentally validate the effectiveness of RRWP features for graph diffusion models and propose extending their proof to additional graph properties that GNNs fail to capture (Xu et al., 2019; Morris et al., 2019). For example, an MLP with input $M^{k}$ with $k=N-1$ for an $N$-node graph can approximate the connected component of each node and the number of vertices in the largest connected component. Additionally, RRWP features can be used to capture cycle-related information.
Appendix E Conditional Generation
@@
where $\gamma$ denotes the guidance weight. In particular, the case with $\gamma=1$ corresponds to standard conditional generation, while $\gamma=0$ represents standard unconditional generation. As $\gamma$ increases, the conditioning effect described by ${\left(\frac{R^{\theta}_{t}(x,\tilde{x}|y)}{R^{\theta}_{t}(x,\tilde{x})}\right)}^{\gamma-1}$ is strengthened, thereby enhancing the quality of the generated samples. We observed $\gamma=2.0$ to be the best performing value for our digital pathology experiments (Sec. G.1), as detailed in Tab. 5.
@@
Appendix F Experimental Details
@@
F.1 Architecture and Additional Features
Denoising Neural Architecture
DeFoG’s denoising neural network takes a noisy graph $G_{t}$ as input and predicts the clean marginal probability for each node $x^{(n)}$ via $p^{\theta,(n)}_{1|t}(\cdot|G_{t})$ and for each edge $e^{(ij)}$ via $p^{\theta,(ij)}_{1|t}(\cdot|G_{t})$. This formulation boils down the graph generative task to a graph-to-graph mapping. While both message-passing layers and graph transformers can be used for this task, graph transformers have empirically outperformed message-passing layers in graph generation (Qin et al., 2023). DeFoG thus adopts the transformer architecture of Vignac et al. (2022), using multi-head attention layers which encode node, edge, and graph-level features while preserving node permutation equivariance.
Enhancing Model Expressivity
Graph neural networks, including graph transformers, have inherent limitations in their expressive power (Xu et al., 2019; Zhu et al., 2023). An usual approach to overcome the limited representation power of graph neural networks consists of explicitly augmenting the inputs with features that the networks would otherwise struggle to learn. We adopt Relative Random Walk Probabilities (RRWP) encodings that are proved to be expressive for both discriminative (Ma et al., 2023) and generative settings (Siraudin et al., 2024). RRWP encodes the likelihood of traversing from one node to another in a graph through random walks of varying lengths, offering insights into graph dynamics across different hop distances. In particular, given a graph with an adjacency matrix $A$, we generate $K-1$ powers of its degree-normalized adjacency matrix, $M=D^{-1}A$, i.e., $\left[I,M,M^{2},\ldots,M^{K-1}\right]$. We concatenate the diagonal entries of each power to their corresponding node embedding, while combining and appending the non-diagonal to their corresponding edge embeddings. In addition to their enhanced expressiveness, RRWP features are easy to compute and thus offer a notably more efficient and scalable alternative to the resource-intensive spectral and cycle features (Vignac et al., 2022) commonly employed in prior works (see Sec. G.4).
F.2 Dataset Details
F.2.1 Synthetic Datasets
Here, we describe the datasets employed in our experiments and outline the specific metrics used to evaluate model performance on each dataset. Additional visualizations of example graphs from each dataset, along with generated graphs, are provided in Figures 23, 25 and 27.
@@ F.2.2 Molecular datasets
Then, we include the Guacamol benchmark (Brown et al., 2019), based on the ChEMBL 24 database (Mendez et al., 2019). This dataset comprises synthesized molecules, tested against biological targets, with sizes ranging from 2 to 88 heavy atoms.
Lastly, we also include the ZINC250k dataset (Sterling & Irwin, 2015), which contains 249,455 molecules with up to 38 heavy atoms from 9 element types. We evaluate DeFoG’s performance under the same setting of previous works (Jo et al., 2024).
@@
For the QM9 dataset, we follow the dataset splits and evaluation metrics outlined by Vignac et al. (2022). For the Moses and Guacamol benchmarks, we adhere to the training setups and evaluation metrics proposed by Polykovskiy et al. (2020) and Brown et al. (2019), respectively. Note that Guacamol includes molecules with charges; therefore, the generated graphs are converted to charged molecules based on the relaxed validity criterion used by Jo et al. (2022) before being translated to their corresponding SMILES representations. The validity, uniqueness, and novelty metrics reported by the Guacamol benchmark are actually V, V.U., and V.U.N., and are referred to directly as V, V.U., and V.U.N. in the table for clarity. For ZINC250k, we adopt the standard evaluation metrics commonly used for this benchmark (see, e.g., Jo et al. (2024); Eijkelboom et al. (2024)), which comprise a subset of the metrics described above.
F.2.3 Digital Pathology Datasets
@@
Graphs, with their natural ability to represent relational data, are widely used to capture spatial biological dependencies in tissue images. This approach has proven successful in digital pathology tasks such as microenvironment classification (Wu et al., 2022), cancer classification (Pati et al., 2022), and decision explainability (Jaume et al., 2020). More recently, graph-based methods have been applied to generative tasks (Madeira et al., 2023), and an open-source dataset was made available by Madeira et al. (2024). This dataset consists of cell graphs where the nodes represent biological cells, categorized into 9 distinct cell types (node classes), and edges model local cell-cell interactions (a single class). For further details, refer to Madeira et al. (2024).
@@
Each cell graph in the dataset can be mapped to a TLS (Tertiary Lymphoid Structure) embedding, denoted as $\kappa=[\kappa_{0},\ldots,\kappa_{5}]\in\mathbb{R}^{6}$, which quantifies its TLS content. A graph $G$ is classified as having low TLS content if $\kappa_{1}(G)<0.05$, and high TLS content if $\kappa_{2}(G)>0.05$. Based on these criteria, the dataset is split into two subsets: high TLS and low TLS. In prior work, TLS generation accuracy was evaluated by training generative models on these subsets separately, and verifying if the generated graphs matched the corresponding TLS content label. We compute TLS accuracy as the average accuracy across both subsets. For DeFoG, we conditionally train it on both subsets simultaneously, as described in Appendix E, and compute TLS accuracy based on whether the generated graphs adhere to the conditioning label. Additionally, we report the V.U.N. metric (valid, unique, novel), similar to what is done for the synthetic datasets (see Sec. F.2.1). A graph is considered valid in this case if it is a connected planar graph, as the graphs in these datasets were constructed using Delaunay triangulation.
F.3 Resources
@@
DeFoG’s memory usage matches existing diffusion models, with quadratic complexity in node number due to complete-graph modeling. Rate matrix overhead is negligible, and RRWP features are more efficient to compute than previous alternatives (Vignac et al., 2022; Xu et al., 2024), as shown in Sec. G.4.
@@
ZINC250k 6 38 14 10000 4.8
F.4 Hyperparameter Tuning
The default hyperparameters for training and sampling for each dataset can be found in the provided code repository. In Tab. 5, we specifically highlight their values for the proposed training and sampling strategies (Sec. 3 and Sec. C.1), and conditional guidance parameter (see Appendix E). As the training process is by far the most computationally costly stage, we aim to minimize changes to the default model training configuration. Nevertheless, we demonstrate the effectiveness of these modifications on certain datasets: